What is the molarity of a solution of H3PO4 if 5.00 moles are dissolved in 2.00 L of solution ?
1
Expert's answer
2017-02-28T07:45:31-0500
Solution
n (H_3 〖PO〗_4) = 5.00 moles
V= 2 L
C=n/V
C=n/V=(5.00 moles)/(2 L)=2.50 mol/L=2.50 M
Answer
The molarity of a solution of H3PO4 must be 2.50 mol/L or 2.50 M.